Evaluate the permutation (36, 19)
Use scientific notation. Round to four decimal places as needed.

Answer by ichigo449(30)

In general notation P(n,k) if k>n then this defined to be 0 and otherwise equal to n!/(n-k)! where x! = the product of all positive integers up to, and including, x.Now this is just plugging numbers: 36!/(36-19)! = 36!/17! . This can be simplified by remembering the definition as all numbers up to 17 will appear in both numerator and denominator and thus cancel. So we are truly left with 36*35*34*33*32*31*30*29*28*27*26*25*24*23*22*21*20*19*18 = 1045843337171591729971200000 which is 1.0459 *10^27. Sorry for the ugly result.